OpenStudy (anonymous):

y = log3(x - 3) + 3 how do I graph this

OpenStudy (anonymous):

reflect the graph \(\large y=3^x \) over the line y=x then translate that result 3 units vertically upwards and 3 units horizontally to the right.

OpenStudy (anonymous):

Rewrite the equation as, y-3 = log3 (x-3)... then the graph is the same as, log3 (x) .. but shifted 3 units to right and 3 units upwards.... [ any graph of y-k=f(x-h) is ... h units right and k units up w.r.t graph of y=f(x)]... now the graph of log3 (x) is as  :Now shift the graph by 3 units to right and 3 units up